### v1

# Open-vocabulary Multimodal Emotion Recognition: Dataset, Metric, and Benchmark

Abstract
Multimodal Emotion Recognition (MER) is an important research topic. This paper advocates for a transformative paradigm in MER. The rationale behind our work is that current approaches often rely on a limited set of basic emotion labels, which do not adequately represent the rich spectrum of human emotions. These traditional and overly simplistic emotion categories fail to capture the inherent complexity and subtlety of human emotional experiences, leading to limited generalizability and practicality. Therefore, we propose a new MER paradigm called Open-vocabulary MER (OV-MER), which encompasses a broader range of emotion labels to reflect the richness of human emotions. This paradigm relaxes the label space, allowing for the prediction of arbitrary numbers and categories of emotions. To support this transition, we provide a comprehensive solution that includes a newly constructed database based on LLM and human collaborative annotations, along with corresponding metrics and a series of benchmarks. We hope this work advances emotion recognition from basic emotions to more nuanced emotions, contributing to the development of emotional AI.
1 Introduction
Research on emotions has a history spanning two centuries. As early as the 19th century, Charles Darwin conducted pioneering research about the evolutionary origins and possible purposes of emotions, explaining the emotional expressions of humans and animals [1]. In 1884, William James revealed the process of emotion generation, noting that stimuli trigger activities in the autonomic nervous system, which in turn produces an emotional experience in the brain [2]. With the rapid development of AI, emotions have garnered increasing attention. [3] highlighted the importance of emotions: The question is not whether intelligent machines can have any emotions, but whether machines can be intelligent without any emotions.
Humans convey emotions through various modalities, which gives rise to the task of Multimodal Emotion Recognition (MER) [4]. Its main goal is to integrate multimodal clues to identify human emotional states. The first step in recognizing emotions is how to model them. Currently, emotion models can be broadly classified into two categories: dimensional models and discrete models. The former uses multi-dimensional space to describe emotions, with the arousal-valence dimension being widely adopted; valence describes the pleasantness of a stimulus, while arousal indicates the intensity of emotion provoked by a stimulus [5]. However, these definitions require specialized psychological knowledge, making them abstract and difficult for the ordinary person to understand. This can lead to discrepancies between different annotators, posing challenges for subsequent applications.
Discrete models align more closely with human understanding of emotions. [6] proposed the basic emotion theory, suggesting that there are six basic emotions: anger, disgust, fear, happiness, sadness, and surprise. This theory is widely used in MER, where researchers typically limit the label space to these basic emotions and use multiple annotators to select the most likely label through majority voting. We refer to this task as One-hot MER (OH-MER). Considering that emotions can be compound, researchers further propose Multi-label MER (ML-MER), allowing each sample to have multiple labels [7]. However, both OH-MER and ML-MER generally have limited label spaces and numbers. Nevertheless, [8] pointed out that humans can express approximately 34,000 different emotions. Therefore, restricting the label space and number will overlook some nuanced emotions.
(a) Task comparison
(b) Label comparison
Figure 1: Comparison. (a) Task Comparison: We compare the differences among three tasks (one-hot MER, multi-label MER, and open-vocabulary MER) across three aspects (label space, label number, and annotation manner). (b) Label Comparison: We provide an example to visualize the one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. We observe that OV labels contain richer emotions. See Appendix B for more examples.
In this paper, we extend traditional MER to Open-vocabulary MER (OV-MER), where we allow the prediction of emotions across any number and categories. Figure 1(a) provides comparisons between different tasks. To facilitate further research, we build an initial dataset, define evaluation metrics, and develop baselines: (1) For the dataset, we propose a human-LLM collaboration strategy. Compared to human-only annotation, our strategy can leverage LLM to enhance the label richness (see Section 2); (2) For evaluation metrics, since there is no fixed label space, the model may predict closely related but differently expressed emotions. For example, the ground truth is joyful and the model predicts happy. To provide more reliable evaluation results, we first group similar emotions and specifically design metrics for this task (see Section 3); (3) For baselines, traditional classifiers rely on fixed label spaces. However, OV-MER does not restrict the label space or number of labels, necessitating the definition of new solutions (see Section 4).
A natural question arises: why is open vocabulary so important for MER? The reason is that OV can generate richer and more nuanced emotional representations, leading to more accurate and reliable MER. As illustrated in Figure 1(b), labeling an emotion solely as surprise is not sufficiently informative—we cannot ascertain whether the surprise is positive or negative, which limits its practicality, such as in human-computer interaction (HCI) applications. In contrast, our OV-MER provides emotions like surprise along with additional descriptors such as nervous and dissatisfied, offering a more comprehensive and insightful understanding of the emotional state. Therefore, OV-MER facilitates the transition from basic to nuanced emotion recognition, advancing the development of emotion AI. In summary, we make the following key contributions:
  • Paradigm. We propose a new paradigm in MER called OV-MER. This paradigm transitions from traditional MER to a framework that enables the prediction of any number and category of emotions, thereby advancing emotion AI toward real-world applicability by capturing the full spectrum of human emotions.
  • Groundwork. We lay the groundwork for OV-MER by constructing datasets, defining evaluation metrics, and proposing effective solutions. Our dataset enhances label richness through a human-LLM collaboration strategy. Meanwhile, we introduce new evaluation metrics that leverage emotional relevance to achieve more reliable results.
  • Benchmark. We establish benchmarks for OV-MER by conducting extensive experiments on various baselines and providing detailed analysis. This task can serve as an important evaluation benchmark for MLLMs, challenging their ability to integrate multimodal clues and capture subtle temporal variations in emotional expression.
2 The OV-MERD Dataset Construction
Although the concept of OV-MER is intuitive and holds great promise, its practical implementation faces significant challenges. The main difficulty lies in the broad and subtle range of human emotions, making comprehensive labeling a complex task. Traditional annotation methods are limited by their predefined emotion categories, which are often insufficient for the needs of OV-MER. Therefore, as shown in Figure 2, we propose a human-LLM collaboration strategy that consists of two steps: CLUE-Multi generation and emotion label extraction. Ultimately, we create a dataset called OV-MERD, which offers a richer set of emotions compared to existing datasets (see Table 1).
2.1 CLUE-Multi Generation
During the annotation process, we observe that descriptions generated through human-LLM collaboration are more detailed. This indicates that when annotations are solely performed by humans, they tend to focus on major clues while neglecting minor ones (see Section 5 for more experimental results). In this section, we provide a detailed overview of our human-LLM collaboration strategy.
Figure 2: Dataset construction. (a) CLUE-Multi Generation: For audio and video, we use ALLM and VLLM to extract initial clues, followed by two rounds of manual checks to eliminate errors and duplicates while adding missing content. Each round involves multiple annotators, with no overlap between annotators in the two rounds. Finally, we merge the checked clues with text to generate CLUE-Multi. (b) Ground-truth OV Label Extraction: There are certain differences in the labels extracted from different languages. To eliminate language influence and achieve consensus labels, we merge these labels and conduct manual checks.
Pre-annotation.
Initially, we attempt to directly annotate visual and acoustic clues. However, the descriptions obtained in this way cannot cover all information. Therefore, we explore using other models for pre-annotation. (1) For video, given the strong visual understanding capabilities of GPT-4V (“gpt-4-vision-preview”), we use it as video LLM (VLLM) to pre-label visual clues. Since GPT-4V only supports image input, we uniformly sample each video and convert it into an image sequence. Considering the relatively short duration of our video, we sample three frames as input. (2) For audio, we use the open-source SALMONN [9] as audio LLM (ALLM) for pre-annotation, because GPT-4V does not support audio input.
Manual Check.
As part of our quality assurance procedures, we perform a detailed examination of the pre-annotated results. For visual clues, GPT-4V may generate hallucinated responses, i.e., clues that do not actually exist. Additionally, there are repeated expressions and some temporal association clues are missing. Therefore, we hire annotators to eliminate errors and duplicates, as well as add missing information. For acoustic clues, ALLM struggles to capture emotion-related acoustic features. Hence, we hire annotators to focus on the speaker’s intonation and other emotional clues. To reduce subjective bias, we conduct two rounds of manual checks, each involving different annotators. These annotators are experts in affective computing and are familiar with the definitions of emotions. Ultimately, these checked clues can accurately reflect the video content. Appendix G provides the annotation guideline and layout of the annotation platform.
CLUE-Multi Generation.
Subsequently, we leverage the semantic reasoning capabilities of LLM to integrate all clues. Specifically, we use GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as the LLM and request it to merge textual, audio, and visual clues to infer emotional states. The output of this process is an emotion-related description, denoted as CLUE-Multi (see Figure 2). Overall, this process reflects the collaboration between humans and LLMs. In Section 5, we compare human-only annotation with this strategy, observing that the latter produces more comprehensive descriptions. In Appendix I, we conduct a further in-depth analysis of CLUE-Multi, demonstrating that it provides a complete representation of emotions, supporting both discrete and dimensional emotion recognition.
2.2 Ground-truth OV Label Extraction
Label Extraction.
After that, we use the LLM to extract emotion labels from CLUE-Multi. This process relies on GPT-3.5, which we request to identify emotional states based on the provided descriptions without restricting the label space. See Appendix D for more details.
Label Filtering.
We further explore the language impact. In Figure 2, we first extract emotion labels from English and Chinese descriptions, resulting in $Y_{EE}$ and $Y_{CC}$. Then, we translate these labels into the other language, yielding $Y_{EC}$ and $Y_{CE}$. After that, we calculate their similarity to assess language influence. In this section, we rely on the Jaccard similarity coefficient to measure the similarity between two sets (see Appendix H). Experimental results indicate that labels extracted from different languages exhibit some differences. For example, the similarity score between $Y_{EE}$ and $Y_{CE}$ is 0.82, which may be due to the varying definitions of emotions in different languages. To eliminate language influence and achieve consensus labels, we merge the labels extracted from both languages and conduct manual checks. These checked labels are regarded as the ground truth.
Table 1: Dataset comparison. “A”, “V”, and “T” are abbreviations for audio, video, and text, respectively. See Appendix F for a comprehensive comparison with existing datasets.
Dataset Modality Annotation Type # Categories # Labels per Sample
MOUD [10] A,V,T Dimensional Emotion 1 1
CMU-MOSI [11] A,V,T Dimensional Emotion 1 1
CMU-MOSEI [12] A,V,T Dimensional Emotion 1 1
CH-SIMS [13] A,V,T Dimensional Emotion 1 1
CH-SIMS v2 [14] A,V,T Dimensional Emotion 1 1
SEMAINE [15] A,V,T Dimensional Emotion 5 1
DFEW [16] A,V,T Discrete Emotion 7 1
MSP-IMPROV [17] A,V,T Discrete Emotion 4 1
IEMOCAP [18] A,V,T Discrete Emotion 10 1
MELD [19] A,V,T Discrete Emotion 7 1
MER2023 [20] A,V,T Discrete Emotion 6 1
MER2024 [21] A,V,T Discrete Emotion 6 1
OV-MERD (Ours) A,V,T Discrete Emotion 248
(arbitrary label)
1$\sim$9, most 2$\sim$4
(arbitrary number)
2.3 OV-MERD Dataset
Finally, we construct a dataset called OV-MERD. This dataset is an extension of MER2023 [20], from which we evenly select samples with further annotations. Table 1 compares OV-MERD with existing datasets. We observe that our OV-MERD dataset contains 248 emotion categories and most samples have 2 to 4 labels, far exceeding those in current datasets. In Appendix E, we observe that OV-MERD encompasses a broader range of emotions, including some that have been rarely discussed in previous research, such as shy, nervous, and grateful.
3 Evaluation Metric
Defining evaluation metrics for OV-MER presents significant challenges: (1) OV-MER supports predicting emotions of any category. Thus, the model may predict closely related but differently expressed emotions. To provide more reliable evaluation results, we first group the emotions based on their similarities. (2) OV-MER allows for the prediction of an arbitrary number of labels. Thus, traditional evaluation metrics designed for a fixed number of labels may not be applicable. In this section, we propose set-based evaluation metrics specifically tailored for OV-MER.
3.1 Grouping
We propose two grouping strategies: one based on GPT and the other based on the emotion wheel (EW) [22]. In the following experiments, we default to using GPT-based grouping.
GPT-based Grouping.
The most direct approach is to use GPT-3.5 to group all labels based on their similarity: Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing emotions with the same meaning. Directly output the results. The output format should be a list containing multiple lists. However, the evaluation results may be affected by the API version. For example, if OpenAI deprecates an old API, the results based on that API will become difficult to reproduce. Additionally, this process is costly; evaluating our OV-MERD dataset costs about \$1, and costs will further increase if the dataset is expanded in the future. Therefore, we attempt to find a replacement for GPT-based grouping.
Figure 3: Emotion wheel W1.
EW-based Grouping.
[…]
Take W1 as an example (see Figure 3). Before calculating the metrics, we define some symbols. We group the labels by their levels from the innermost to the outermost as $L_{w_{1}}^{1}$, $L_{w_{1}}^{2}$, and $L_{w_{1}}^{3}$. Specifically, $L_{w_{1}}^{1}=\{\text{mad},\cdots,\text{scared}\}$. Next, we define a function $m_{w_{1}}^{i\rightarrow j}(\cdot)$ that maps the labels in $L_{w_{1}}^{i}$ to the corresponding labels in $L_{w_{1}}^{j}$. From inner to outer ($i<j$), $m_{w_{1}}^{i\rightarrow j}(\cdot)$ is a many-to-one mapping; from outer to inner ($i>j$), $m_{w_{1}}^{i\rightarrow j}(\cdot)$ is a one-to-many mapping. For example, $m_{w_{1}}^{2\rightarrow 1}(\mbox{angry})=\{\mbox{mad}\}$ and $m_{w_{1}}^{1\rightarrow 2}(\mbox{mad})=\{\mbox{offended},\cdots,\mbox{humiliated}\}$. We collect all the labels from these emotion wheels and represent them as EW, i.e., $\{L_{w_{i}}^{j},1\leq i\leq 5,1\leq j\leq 3\}$. We denote the labels in EW as $y_{w}$.
Considering that the emotional categories in EW are still limited, we perform some label expansion operations. Specifically, we repeatedly call GPT-3.5, asking it to generate synonyms for each label. The prompt used is as follows: Please retrieve the synonyms for the following words and output them in a table format. Then, we generate EW-S, i.e., $\{f(y_{w})=\{y_{f}^{1},...,y_{f}^{n}\},y_{w}\in\mbox{EW}\}$, where $f(\cdot)$ is a function that maps each label $y_{w}$ to its synonym $y_{f}$. We also define its inverse function $f^{\prime}(\cdot)$, which maps different synonyms $y_{f}$ back to their base label $y_{w}$.
[…]
(1) M1. We use $g^{\prime}(\cdot)$ to map all labels to their corresponding $y_{f}$.
(2) M2. We use $f^{\prime}(g^{\prime}(\cdot))$ to map all labels to their corresponding $y_{w}$.
(3) M3. We use the emotion wheel during metric calculation. Specifically, we first use $f^{\prime}(g^{\prime}(\cdot))$ to map all labels to their corresponding $y_{w}$. Then, we define two grouping functions, L1 and L2. For L1, we map all labels to their corresponding $L_{w_{i}}^{1}$:
$\begin{cases}y_{w},&\text{if}\;y_{w}\in L_{w_{i}}^{1}\\ m_{w_{i}}^{2\rightarrow 1}(y_{w}),&\text{if}\;y_{w}\in L_{w_{i}}^{2}\\ m_{w_{i}}^{2\rightarrow 1}(m_{w_{i}}^{3\rightarrow 2}(y_{w})),&\text{if}\;y_{w}\in L_{w_{i}}^{3}\\ \end{cases}$ (1)
For L2, we map all labels to their corresponding $L_{w_{i}}^{2}$:
$\begin{cases}\text{select one label in}\;m_{w_{i}}^{1\rightarrow 2}(y_{w}),&\text{if}\;y_{w}\in L_{w_{i}}^{1}\\ y_{w},&\text{if}\;y_{w}\in L_{w_{i}}^{2}\\ m_{w_{i}}^{3\rightarrow 2}(y_{w}),&\text{if}\;y_{w}\in L_{w_{i}}^{3}\\ \end{cases}$ (2)
3.2 Metric Definition
Then, we convert the above emotion grouping information into a function $G(\cdot)$, which can map each label to its group ID. Specifically, suppose $\{y_{i}\}_{i=1}^{M}$ and $\{\hat{y}_{i}\}_{i=1}^{N}$ are the ground truth and predictions, where $M$ and $N$ are the number of labels. We first map each label into its group ID:
$\mathcal{Y}=\{G(x)|x\in\{y_{i}\}_{i=1}^{M}\},\hat{\mathcal{Y}}=\{G(x)|x\in\{\hat{y}_{i}\}_{i=1}^{N}\}.$ (3)
Finally, we design set-based metrics for performance evaluation. Specifically, $\mbox{Precision}_{\mbox{s}}$ indicates the number of correctly predicted labels; $\mbox{Recall}_{\mbox{s}}$ indicates whether the prediction covers all ground truth; and $\mbox{F}_{\mbox{s}}$ is the harmonic mean of two metrics, which is used for the final ranking:
$\mbox{Precision}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\hat{\mathcal{Y}}|},\;\mbox{Recall}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\mathcal{Y}|},\;\mbox{F}_{\mbox{s}}=2\times\frac{\mbox{Precision}_{\mbox{s}}\times\mbox{Recall}_{\mbox{s}}}{\mbox{Precision}_{\mbox{s}}+\mbox{Recall}_{\mbox{s}}}.$ (4)
4 Baselines for OV-MER
4.1 CLUE Generation
Figure 2 illustrates the generation process of CLUE-Multi, where we combine text with manually checked visual and acoustic clues. In this section, we further introduce the following variants.
CLUE-A/T/V.
To reveal the modality impact, we propose three variants of CLUE-Multi: CLUE-Audio, CLUE-Text, and CLUE-Video. Their generation process is illustrated in Figure 4. Specifically, since audio contains both acoustic information (e.g., the speaker’s intonation) and content, we merge the checked acoustic clues and text to generate CLUE-Audio; for CLUE-Text, we only use the text to infer emotional states; since the visual content does not contain audio and text, for CLUE-Video, we only use the checked visual clues. See Appendix J for more examples.
CLUE-MLLM.
MLLMs can address various multimodal tasks, with the main idea being to align information from other modalities with LLMs. Since emotion recognition relies on temporal information, we choose models that support at least video or audio. To generate CLUE-MLLM, we first use ALLM or VLLM to extract emotion-related descriptions, and then combine these descriptions with text using LLM. Compared with CLUE-Multi, this process does not use manually checked clues. Appendix K provides the model card and relevant prompts. All models are implemented in PyTorch, and all inference processes are carried out using a 32G NVIDIA Tesla V100 GPU.
Figure 4: Baselines. (a) Preliminary: We begin by defining some preliminary symbols. (b) CLUE Generation: CLUE-Video and CLUE-Audio use manually-checked clues; CLUE-Text relies solely on text; CLUE-MLLM does not involve manual checks and directly uses the outputs from ALLM or VLLM. (c) Metric Calculation: We rely on CLUE to predict emotion labels. Due to variations in labels extracted from different languages, we report results across different languages.
4.2 Metric Calculation
As shown in Figure 2, there are certain differences in the labels extracted from different languages. Therefore, we report the results for both English and Chinese descriptions. In Figure 4, for the Chinese branch, we first extract OV labels and then translate them into English; for the English branch, we directly extract OV labels. Finally, we compute the evaluation metrics with the ground truth. It is worth noting that the OV labels extracted from the monolingual CLUE-Multi differ from the ground truth. Our ground truth combines the labels extracted from different languages and undergoes further manual checks (see Figure 2).
Table 2: Baseline results. Figure 4 shows the metric calculation process.
Model L V A English Chinese
$\mbox{F}_{\mbox{s}}$uparrow $\mbox{Precision}_{\mbox{s}}$uparrow $\mbox{Recall}_{\mbox{s}}$uparrow $\mbox{F}_{\mbox{s}}$uparrow $\mbox{Precision}_{\mbox{s}}$uparrow $\mbox{Recall}_{\mbox{s}}$uparrow
Heuristic Baseline
Random $\times$ $\times$ $\times$ 17.42±0.01 24.85±0.15 13.42±0.04 16.59±0.00 24.70±0.00 12.48±0.00
CLUE-MLLM
Qwen-Audio $\surd$ $\times$ $\surd$ 38.13±0.05 49.42±0.18 31.04±0.00 41.14±0.07 53.71±0.00 33.34±0.09
OneLLM $\surd$ $\times$ $\surd$ 42.84±0.06 45.92±0.05 40.15±0.06 46.17±0.02 52.07±0.06 41.47±0.08
Otter $\surd$ $\surd$ $\times$ 43.51±0.09 50.71±0.10 38.09±0.09 46.22±0.01 52.65±0.16 41.18±0.08
Video-LLaMA $\surd$ $\surd$ $\times$ 44.73±0.14 44.14±0.13 45.34±0.15 47.26±0.03 47.98±0.07 46.56±0.01
VideoChat $\surd$ $\surd$ $\times$ 45.53±0.11 42.90±0.27 48.49±0.10 45.57±0.03 47.20±0.12 44.05±0.05
SECap $\surd$ $\times$ $\surd$ 45.72±0.09 54.52±0.15 39.37±0.05 45.57±0.13 55.55±0.23 38.64±0.08
PandaGPT $\surd$ $\surd$ $\surd$ 45.89±0.20 50.03±0.01 42.38±0.33 47.33±0.04 53.01±0.08 42.75±0.11
Video-LLaVA $\surd$ $\surd$ $\times$ 47.07±0.16 48.58±0.02 45.66±0.29 49.21±0.06 53.95±0.03 45.23±0.13
SALMONN $\surd$ $\times$ $\surd$ 47.96±0.04 50.20±0.04 45.92±0.04 48.24±0.03 52.24±0.00 44.82±0.05
VideoChat2 $\surd$ $\surd$ $\times$ 49.07±0.26 54.72±0.41 44.47±0.15 48.86±0.05 57.12±0.08 42.68±0.04
Video-ChatGPT $\surd$ $\surd$ $\times$ 50.52±0.06 54.03±0.04 47.44±0.07 54.73±0.00 61.15±0.10 49.52±0.06
OneLLM $\surd$ $\surd$ $\times$ 50.52±0.07 55.93±0.09 46.06±0.06 51.44±0.08 56.43±0.04 47.26±0.11
LLaMA-VID $\surd$ $\surd$ $\times$ 51.25±0.09 52.71±0.18 49.87±0.00 52.01±0.02 57.30±0.00 47.61±0.03
mPLUG-Owl $\surd$ $\surd$ $\times$ 52.73±0.13 54.54±0.13 51.04±0.13 50.95±0.06 56.40±0.11 46.47±0.18
Chat-UniVi $\surd$ $\surd$ $\times$ 53.08±0.01 53.68±0.00 52.50±0.02 53.86±0.02 58.54±0.01 49.86±0.03
GPT-4V $\surd$ $\surd$ $\times$ 55.51±0.05 48.52±0.07 64.86±0.00 57.21±0.01 54.61±0.02 60.07±0.01
CLUE-M/A/T/V
CLUE-Text $\surd$ $\times$ $\times$ 46.00±0.06 54.41±0.15 39.84±0.01 43.11±0.25 50.69±0.26 37.50±0.23
CLUE-Video $\times$ $\surd$ $\times$ 60.55±0.13 63.29±0.08 58.05±0.16 61.73±0.10 66.47±0.13 57.62±0.08
CLUE-Audio $\surd$ $\times$ $\surd$ 65.35±0.04 67.54±0.08 63.30±0.00 68.56±0.07 70.10±0.06 67.07±0.08
CLUE-Multi $\surd$ $\surd$ $\surd$ 80.05±0.24 80.03±0.37 80.07±0.10 85.16±0.03 87.09±0.00 83.31±0.05
5 Results and Discussion
In this section, we report the experimental results and provide an in-depth analysis. We default to using GPT-based grouping and employ GPT-3.5 (“gpt-3.5-turbo-16k-0613”) as LLM. We generally report evaluation results in both languages, but if no specific language is mentioned, we default to reporting results for the English branch. To mitigate the impact of randomness, we conduct each experiment twice and report the average scores and standard deviations.
Main Results on CLUE-M/A/T/V.
Table 2 shows the performance of CLUE-M/A/T/V. We observe that CLUE-Multi performs the best. This is because emotions are conveyed through various modalities, and integrating multimodal clues can better understand human emotions. In contrast, CLUE-Text performs the worst. This is because, in MER, text is emotionally ambiguous; without integrating other modalities, we cannot accurately understand the emotions conveyed through text.
Main Results on CLUE-MLLM.
Table 2 presents the results of CLUE-MLLM. Additionally, we introduce a heuristic baseline called Random, where we randomly select a label from basic emotions. This baseline reflects the lower bound of performance. We observe that MLLM generally outperforms Random, indicating that MLLM can partially address OV-MER. However, the performance of MLLM remains unsatisfactory, highlighting the limitations of existing MLLMs and the challenges of OV-MER. Furthermore, models that perform well in Chinese often perform well in English. These results suggest that the impact of language differences on rankings is limited.
Table 3: Performance of combinations of different MLLMs.
Model English Chinese
$\mbox{F}_{\mbox{s}}\uparrow$ $\mbox{Precision}_{\mbox{s}}\uparrow$ $\mbox{Recall}_{\mbox{s}}\uparrow$ $\mbox{F}_{\mbox{s}}\uparrow$ $\mbox{Precision}_{\mbox{s}}\uparrow$ $\mbox{Recall}_{\mbox{s}}\uparrow$
Audio + Text
SECap 45.72±0.09 54.52±0.15 39.37±0.05 45.57±0.13 55.55±0.23 38.64±0.08
SALMONN 47.96±0.04 50.20±0.04 45.92±0.04 48.24±0.03 52.24±0.00 44.82±0.05
Video + Text
Video-ChatGPT 50.52±0.06 54.03±0.04 47.44±0.07 54.73±0.00 61.15±0.10 49.52±0.06
mPLUG-Owl 52.73±0.13 54.54±0.13 51.04±0.13 50.95±0.06 56.40±0.11 46.47±0.18
Chat-UniVi 53.08±0.01 53.68±0.00 52.50±0.02 53.86±0.02 58.54±0.01 49.86±0.03
Audio + Video + Text
SECap + mPLUG-Owl 56.69±0.03 50.05±0.23 65.38±0.33 54.99±0.23 51.65±0.27 58.79±0.16
SECap + Video-ChatGPT 56.90±0.08 52.03±0.04 62.79±0.14 56.49±0.02 56.50±0.01 56.48±0.05
SECap + Chat-UniVi 57.34±0.16 48.85±0.29 69.41±0.13 56.19±0.13 52.38±0.07 60.59±0.22
SALMONN + Video-ChatGPT 58.19±0.23 53.16±0.17 64.26±0.31 55.05±0.16 53.44±0.14 56.76±0.19
SALMONN + Chat-UniVi 58.43±0.06 51.62±0.00 67.31±0.15 56.93±0.06 51.65±0.06 63.42±0.06
SALMONN + mPLUG-Owl 58.70±0.04 51.77±0.01 67.76±0.11 55.62±0.21 51.74±0.19 60.14±0.23
(a) Clue length (H)
(b) Label count (H)
(c) Word cloud (H)
(d) Clue length (H+L)
(e) Label count (H+L)
(f) Word cloud (H+L)
Figure 5: Human-only annotation (H) vs. Human-LLM (H+L) collaboration.
(a) PandaGPT
(b) Video-ChatGPT
(c) Video-LLaMA
(d) VideoChat
(e) VideoChat2
(f) mPLUG-Owl
(g) Qwen-Audio
(h) Video-LLaVA
(i) LLaMA-VID
(j) Chat-UniVi
Figure 6: Performance comparison of different strategies for generating CLUE-MLLM.
Effectiveness of Multimodal Fusion.
In Table 3, we select the best-performing ALLMs and VLLMs and explore whether their combinations can lead to better performance. We observe that these combinations bring performance improvements, with some even surpassing GPT-4V. Therefore, we can more effectively address OV-MER by leveraging the strengths of different MLLMs.
Figure 7: Different ways for generating CLUE-MLLM.
Human-only vs. Human-LLM Collaboration.
This paper proposes a human-LLM collaborative strategy. In this section, we discuss the necessity of this strategy and test the performance of human-only annotation. In Figure 5, we compare different strategies from three aspects: the length distribution of generated clues, the distribution of label counts, and the word cloud. We observe that the descriptions generated through human-LLM collaboration are longer, indicating that human-only annotations often focus on primary clues while neglecting minor ones. With the pre-annotation and semantic reasoning capabilities of LLMs, we can generate richer and more comprehensive descriptions. Additionally, human-LLM collaboration can produce a broader range of emotions and provide more diverse labels for each sample. These results validate the effectiveness of our strategy.
Ablation Study on CLUE-MLLM.
This section reveals the impact of different CLUE-MLLM generation strategies. Figure 7 introduces three methods: 1) S0 does not use text and inputs the video into MLLM; 2) S1 inputs both text and video into MLLM simultaneously; 3) S2 first uses MLLM to extract descriptions and then combines with text using another LLM, consistent with the strategy shown in Figure 4. In Figure 6, we present the performance of different strategies. Generally, S1 and S2 outperform S0, indicating the importance of text content in OV-MER. Moreover, S2 typically performs better than S1. The reason is that inputting both video and text into the MLLM simultaneously increases the task difficulty, and current MLLMs may struggle to handle complex prompts. S2 divides this process into two steps, reducing task complexity and achieving better performance. Therefore, we adopt S2 as the default strategy.
GPT-based vs. Matching-based Metrics.
In Table 2, CLUE-Multi performs the best. This leads to a hypothesis: Do sentences that are more “similar” to CLUE-Multi yield better emotion recognition performance? The most common way to measure “similarity” is through matching-based metrics, with BLEU11{1}1, BLEU44{4}4, METEOR, and ROUGE1 being the most widely used. Therefore, we use CLUE-MLLM as input and calculate both GPT-based and matching-based metrics (see Figure 8(a)), followed by calculating their PCC scores (see Figure 8(b)). From the experimental results, we have some interesting observations. First, the same metric across different languages typically exhibits high correlations. For example, “F(C)” and “F(E)” have a high PCC score. However, the correlation between GPT-based and matching-based metrics is not strong. For example, the highest PCC score between “F(E)” and matching-based metrics is 0.77. Therefore, a high score in matching-based metrics does not necessarily indicate good performance in OV-MER.
GPT-based vs. EW-based Grouping.
This paper proposes two grouping strategies: GPT-based grouping and EW-based grouping. In this section, we explore the relationship between them and analyze whether the EW-based method can replace the GPT-based method. We calculate the scores for different types of EW-based grouping strategies, with the average score denoted as M-avg. Table 4 reports $\mbox{F}_{\mbox{s}}$, as this metric is used for the final ranking, and we also compute the PCC scores between different metrics. We observe that the PCC score between M-avg and GPT is relatively high, indicating that the EW-based method can replace GPT.
(a) Metric Calculation
(b) Correlation Analysis
Figure 8: GPT- vs. Matching-based metrics.
Table 4: GPT-based vs. EW-based grouping. We calculate the PCC score to reveal their correlation.
Model GPT M1 M2 M3-W1 M3-W2
L1 L2 L1 L2
Qwen-Audio 38.13±0.05 20.49±0.01 23.37±0.01 43.85±0.03 26.60±0.01 41.52±0.28 26.68±0.01
Otter 43.51±0.09 22.21±0.06 27.99±0.02 49.75±0.11 33.50±0.06 49.93±0.11 33.04±0.06
Video-LLaMA 44.73±0.14 23.56±0.08 28.39±0.17 52.90±0.12 36.08±0.13 53.60±0.04 35.33±0.08
VideoChat 45.53±0.11 22.15±0.00 26.24±0.08 47.79±0.07 32.64±0.07 47.76±0.11 32.14±0.04
SECap 45.72±0.09 26.52±0.01 32.88±0.03 52.26±0.03 37.55±0.03 52.11±0.03 37.71±0.03
Video-LLaVA 47.07±0.16 25.47±0.12 30.73±0.11 54.65±0.10 37.65±0.24 54.54±0.02 38.25±0.22
SALMONN 47.96±0.04 23.57±0.02 28.83±0.03 54.90±0.15 38.93±0.15 54.29±0.06 37.79±0.07
VideoChat2 49.07±0.26 26.92±0.09 31.40±0.10 52.38±0.13 36.44±0.11 53.56±0.13 36.91±0.11
Video-ChatGPT 50.52±0.06 28.99±0.04 34.05±0.05 57.66±0.04 41.48±0.09 57.37±0.00 40.95±0.08
LLaMA-VID 51.25±0.09 28.28±0.04 32.85±0.03 56.59±0.04 41.22±0.02 57.49±0.03 40.39±0.04
mPLUG-Owl 52.73±0.13 27.47±0.17 32.47±0.19 57.60±0.23 41.32±0.04 56.32±0.26 40.83±0.07
Chat-UniVi 53.08±0.01 28.89±0.02 33.23±0.08 57.00±0.06 42.25±0.04 57.50±0.03 42.43±0.03
PCC score 0.887 0.857 0.911 0.940 0.913 0.942
Model M3-W3 M3-W4 M3-W5 M-avg
L1 L2 L1 L2 L1 L2
Qwen-Audio 39.46±0.28 30.65±0.01 36.64±0.03 27.33±0.01 35.89±0.08 29.66±0.01 31.84
Otter 51.03±0.04 37.12±0.00 47.54±0.00 34.77±0.00 50.51±0.03 35.54±0.00 39.41
Video-LLaMA 47.50±0.20 36.50±0.25 52.97±0.09 35.78±0.14 46.39±0.12 34.77±0.23 40.31
VideoChat 46.78±0.11 34.37±0.03 49.53±0.15 32.82±0.01 45.93±0.18 32.85±0.04 37.58
SECap 50.77±0.03 40.49±0.03 50.43±0.03 38.21±0.03 49.97±0.03 40.25±0.03 42.43
Video-LLaVA 52.29±0.05 40.58±0.15 52.45±0.06 39.91±0.13 52.97±0.10 39.69±0.10 43.27
SALMONN 56.25±0.01 43.01±0.02 50.53±0.09 38.54±0.03 53.65±0.04 42.09±0.02 43.53
VideoChat2 52.14±0.23 40.57±0.14 50.63±0.19 39.64±0.18 51.37±0.14 39.89±0.15 42.65
Video-ChatGPT 55.50±0.13 44.15±0.18 55.24±0.02 42.42±0.05 52.93±0.05 41.54±0.14 46.02
LLaMA-VID 55.12±0.05 44.06±0.01 56.62±0.15 42.42±0.03 53.03±0.08 41.65±0.04 45.81
mPLUG-Owl 55.67±0.19 43.71±0.13 55.06±0.17 40.67±0.19 54.44±0.13 42.00±0.18 45.63
Chat-UniVi 56.80±0.01 45.66±0.05 55.86±0.07 41.97±0.09 55.81±0.02 43.61±0.05 46.75
PCC score 0.904 0.927 0.899 0.922 0.885 0.894 0.942
6 Discussion
Limitations.
The main contribution of this paper is the definition of a new task and the conduct of foundational research, including the development of datasets, metrics, and solutions. In the future, we plan to design more effective frameworks to solve OV-MER. Meanwhile, we have evaluated some representative MLLMs, but not all models are covered. In the future, we will broaden the scope of our evaluations. In addition, although the OV-MERD dataset is currently small in scale, we intend to expand its size in the future. In fact, preliminary experiments on the OV-MERD dataset provide guidance for our further improvements.
Label Importance.
In OV-MER, we do not assign different levels of importance to each label. We believe that every emotion holds equal significance, and neglecting anyone can impact the performance of downstream tasks. For example, if a human-computer interaction system only captures basic emotions while overlooking nuanced ones, it may fail to generate appropriate responses.
7 Conclusion
This paper extends traditional MER to OV-MER, allowing for the prediction of arbitrary numbers and types of emotions. To facilitate further research, we construct an initial dataset, define evaluation metrics, and propose solutions. We observe that current MLLMs struggle to achieve satisfactory results, as this task requires consideration of multimodal clues and subtle temporal changes, placing higher demands on MLLMs. Additionally, EW-based metrics can replace GPT-based metrics, thus reducing evaluation costs while ensuring reproducibility. This paper advances current research from basic to nuanced emotion recognition, which is crucial for building emotion AI.
Ethics Statement
The raw data of the OV-MERD dataset comes from the MER2023 dataset, from which we evenly select some samples with further annotation. Therefore, we do not collect new data; we just re-annotate existing data. This annotation process has received consent from the dataset owners and has passed our internal review. During the annotation process, we generously pay each annotator approximately ¥3,000 (around \$280), which is considered high. After proofreading our annotation results, we find that the annotations focus on the multimodal clues present in the videos, without any discriminatory annotations. Additionally, we restrict the use of the OV-MERD dataset to non-commercial purposes under the CC BY-NC 4.0 license. This license clearly outlines the correct and responsible use of our dataset. Details of our license are provided in the supplementary materials. In summary, this paper does not involve any ethical issues.
Reproducibility Statement
This paper provides the source code and intermediate results in the supplementary materials. Since we build a dataset for OV-MER, we also offer a complete description of the data construction process in Section 2 and Appendix. In summary, we have made every effort to ensure its reproducibility.
Appendix A Related Work
Multimodal Emotion Recognition.
MER has rapidly developed in recent years [23]. Current research mainly focuses on building more efficient architectures to achieve higher accuracy on benchmark datasets [24]. For example, [11] proposed a tensor fusion network that addressed the MER task by leveraging interactions among unimodal, bimodal, and trimodal inputs. [25] introduced a Transformer-based model that learned implicit alignment between different modalities and achieved promising results. [4] further established MERBench, involving various features, fusion strategies, and datasets. In emotion recognition, benchmark datasets usually limit the label space to basic emotions and use majority voting to determine the most likely one or more labels [20, 7]. However, emotional categories extend far beyond basic emotions. Restricting the label space will inevitably overlook some nuanced emotions. To address this issue, we extend traditional MER to OV-MER, which allows for the prediction of any number and categories of emotions.
Open Vocabulary Learning.
Its main goal is to identify categories beyond the annotated label space [26], which has been applied in various fields, such as object detection [27], segmentation [28], and scene understanding [29]. For example, the object detection dataset COCO [30] contains 80 categories, while objects in the real world are nearly infinite, highlighting the importance of open vocabulary learning. This paper makes the first attempt to address MER in an open-vocabulary manner. Compared to other tasks, MER is more difficult as it requires considering multimodal clues and subtle temporal variations.
Appendix B More Examples
In Figures 9$\sim$11, we provide more examples to visualize the difference between one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. We use DemoAI to remove personal information.
Figure 9: Example1.
Figure 10: Example2.
Figure 11: Example3.
Appendix C Summary of Abbreviations
In Table 5, we summarize the main abbreviations and their meanings.
Table 5: Summary of main abbreviations.
Category Abbreviation Explanation
Label OH label One-hot label, the most likely label in a limited set of basic emotions.
OV labels Open-vocabulary labels, a set of labels in an unlimited label space.
Task MER Multimodal Emotion Recognition, which aims to recognize the one-hot emotion label.
OV-MER Open-vocabulary MER, which aims to identify the OV emotion labels.
Dataset OV-MERD This is a dataset we built for the OV-MER task.
Metric EW Emotion Wheel.
M1, M2, M3 Different grouping strategies based on the emotion wheel.
$\mbox{Precision}_{\mbox{s}}$, $\mbox{Recall}_{\mbox{s}}$, $\mbox{F}_{\mbox{s}}$ Metrics defined for OV-MER.
Model LLM Large Language Model. Large-scale models and only process text.
ALLM Audio LLM. Different from LLM, it can also process audio input.
VLLM Video LLM. Different from LLM, it can also process video input.
MLLM Multimodal LLM. Unlike LLM, it can process at least one more modality (e.g., audio or video). Thus, MLLM includes ALLM and VLLM.
Description CLUE-Multi It uses the checked acoustic and visual clues to generate descriptions.
CLUE-Audio Different from CLUE-Multi, it only uses checked acoustic clues.
CLUE-Video Different from CLUE-Multi, it only uses checked visual clues.
CLUE-Text It only relies on text to generate descriptions.
CLUE-A/T/V Any of CLUE-Audio, CLUE-Text, and CLUE-Video.
CLUE-M/A/T/V Any of CLUE-Multi, CLUE-Audio, CLUE-Text, and CLUE-Video.
CLUE-MLLM It uses the output from MLLM without any manual checking process.
S0, S1, S2 Different CLUE-MLLM generation strategies.
Appendix D Details in Dataset Construction
Figure 2 presents our dataset construction process. In Table 6, we provide prompts and corresponding models used in this process. In Figure 12, we visualize the output of the main steps.
Table 6: Prompts and corresponding models used in the dataset construction process.
Function (Model) Prompt
#1 Pre-label visual clue (VLLM) As an expert in the field of emotions, please focus on facial expressions, body language, environmental cues, and events in the video and predict the emotional state of the character. Please ignore the character’s identity. We uniformly sample 3 frames from this video. Please consider the temporal relationship between these frames and provide a complete description of this video. Avoid using descriptions like “the first image” and “the second image”, and instead use terms like “beginning”, “middle”, and “end” to denote the progression of time.
#2 Pre-label acoustic clue (ALLM) As an expert in the field of emotions, please focus on the acoustic information in the audio to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual.
#3 Merge (LLM) Please act as an expert in the field of emotions. We provide acoustic and visual clues that may be related to the character’s emotional state, along with the original subtitle of the video. Please analyze which parts can infer the emotional state and explain the reasons. During the analysis, please integrate the textual, audio, and visual clues.
[…]
#5 OV label extraction (LLM) Please assume the role of an expert in the field of emotions. We provide clues that may be related to the emotions of the characters. Based on the provided clues, please identify the emotional states of the main characters. Please separate different emotional categories with commas and output only the clearly identifiable emotional categories in a list format. If none are identified, please output an empty list.
Figure 12: An example to visualize the output of the main steps.
Appendix E One-hot vs. OV Labels
This section provides a deeper comparison between the one-hot labels in the MER2023 dataset and the OV labels in the OV-MERD dataset. Figure 13 shows the word cloud and label number distribution of OV labels. In Figure 13(a), we observe that OV labels cover a wider variety of emotions, some of which (such as shy, nervous, and grateful) are rarely discussed in previous datasets. In Figure 13(b), we notice that most samples have about 2 to 4 labels, much more than the traditional task where each sample is assigned only one emotion. Therefore, OV-MER provides richer labels.
(a) Word cloud
(b) Label number distribution
Figure 13: Word cloud and label number distribution of OV labels.
In Table 7, we report the performance of one-hot labels in OV-MER. We observe that one-hot labels have high $\mbox{precision}_{\mbox{s}}$ but low $\mbox{recall}_{\mbox{s}}$, indicating that one-hot labels are correct but not comprehensive. Due to the limited label space and the constrained number of labels, one-hot labels cannot cover all emotions, highlighting the limitations of traditional MER and the importance of OV-MER. Additionally, these results reflect the necessity to use $\mbox{F}_{\mbox{s}}$ for the final ranking, which can balance accuracy and completeness.
Table 7: Performance of one-hot labels in OV-MER.
Language $\mbox{F}_{\mbox{s}}\uparrow$ $\mbox{Precision}_{\mbox{s}}\uparrow$ $\mbox{Recall}_{\mbox{s}}\uparrow$
English 65.71±0.06 92.17±0.00 51.05±0.08
Chinese 66.16±0.02 93.07±0.00 51.32±0.03
Figure 14 shows the emotion distribution of OV labels. We observe that the number of samples for different emotions follows a long-tail distribution. These results indicate that OV labels not only cover some major labels but also capture subtle emotions that occur infrequently.
Figure 14: Emotion distribution of OV labels.
Appendix F Dataset Comparison
This paper introduces a new task, OV-MER, and constructs a dataset for this task called OV-MERD. Table 8 compares OV-MERD with existing datasets. The annotation types of these datasets can be broadly categorized into two types: dimensional emotions and discrete emotions. We classify sentiment analysis datasets (e.g., CMU-MOSI) as dimensional datasets because the definition of sentiment intensity overlaps with the valence in dimensional emotions. We observe that OV-MERD contains 248 emotion categories, with most samples having 2 to 4 labels, significantly exceeding the number of labels in existing datasets. In the future, as the scale of the dataset increases, the number of candidate labels can be further expanded.
Table 8: Dataset comparison. “I”, “A”, “V”, and “T” are abbreviations for image, audio, video, and text, respectively.
Dataset Modality Annotation Type # Categories # Labels per Sample
SST [31] T Dimensional 1 1
Cornell [32] T Dimensional 1 1
Large Movie [33] T Dimensional 1 1
ICT-MMMO [34] A,V,T Dimensional 1 1
YouTube [35] A,V,T Dimensional 1 1
MOUD [10] A,V,T Dimensional 1 1
CMU-MOSI [11] A,V,T Dimensional 1 1
CMU-MOSEI [12] A,V,T Dimensional 1 1
CH-SIMS [13] A,V,T Dimensional 1 1
CH-SIMS v2 [14] A,V,T Dimensional 1 1
VAM [36] A,V,T Dimensional 3 1
SEMAINE [15] A,V,T Dimensional 5 1
AFEW-VA [37] A,V,T Dimensional 2 1
SEWA[38] A,V,T Dimensional 3 1
SFEW 2.0 [39] I Discrete 7 1
FER-2013 [40] I Discrete 7 1
EmotioNet [41] I Discrete 23 1
RAF-DB [7] I Discrete 19 1$\sim$2
AffectNet [42] I Discrete 7 1
ExpW [43] I Discrete 7 1
AFEW 7.0 [44] A,V,T Discrete 7 1
MAFW [45] A,V,T Discrete 11 1
DFEW [16] A,V,T Discrete 7 1
CREMA-D [46] A,V,T Discrete 6 1
MSP-IMPROV [17] A,V,T Discrete 4 1
RAVDESS [47] A,V,T Discrete 8 1
IEMOCAP [18] A,V,T Discrete 10 1
MELD [19] A,V,T Discrete 7 1
MER2023 [20] A,V,T Discrete 6 1
MER2024 [21] A,V,T Discrete 6 1
OV-MERD (Ours) A,V,T Discrete 248
(arbitrary label)
1$\sim$9, most 2$\sim$4
(arbitrary number)
Appendix G Annotation Details
This section presents our annotation guidelines and the layout of the annotation platform. Our annotation process relies on the Label Studio [48] toolkit. As shown in Figure 2, there are two parts that require manual checking: 1) the pre-annotated acoustic and visual clues; 2) the merged open-vocabulary labels. To reduce subjective bias, we hire eight annotators who are experts in affective computing and familiar with the definitions of emotions. Additionally, we conduct two rounds of checks with no overlap among annotators in each round. Specifically, in the first round, we randomly select four annotators to check the clues and labels; in the second round, we merge the clues and labels reviewed by the first four annotators and ask another four annotators to perform a second round of checks. Ultimately, we find that these checked clues and labels are well-aligned with the video content.
Figure 15 shows the layout of the annotation platform used for manually checking acoustic and visual clues. During annotation, we use the following instructions: We provide pre-labeled acoustic and visual clues. Please manually check these clues, remove errors, and add missing information. To reduce the difficulty of annotation, we design an interface with a time slider, allowing annotators to start playing the video from any frame to capture subtle changes within the video.
Figure 15: Layout of the annotation platform used for manually checking acoustic and visual clues.
Figure 16 displays the layout of the annotation platform used for manually checking emotional labels. During annotation, we use the following instructions: Please select all labels that match the character’s emotional state in the “candidate emotions”. If the provided candidate labels cannot perfectly describe the character’s emotional state, you can also manually add new labels to the “other emotions” part. Specifically, annotators need to label two parts. First, we list all candidate labels from which annotators can choose what they believe to be the correct labels; second, when the candidate labels cannot perfectly describe the emotions, annotators can manually add additional labels in the “other emotions” part.
Figure 16: Layout of the annotation platform used for manually checking emotional labels.
Appendix H Jaccard Similarity Coefficient
Figure 2 uses the Jaccard similarity coefficient to measure the similarity between two sets, which is slightly different from the evaluation metrics defined in Section 3. Specifically, in Section 3, we use the following metrics:
$\mbox{Precision}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\hat{\mathcal{Y}}|},\;\mbox{Recall}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\mathcal{Y}|},\;\mbox{F}_{\mbox{s}}=2\times\frac{\mbox{Precision}_{\mbox{s}}\times\mbox{Recall}_{\mbox{s}}}{\mbox{Precision}_{\mbox{s}}+\mbox{Recall}_{\mbox{s}}}.$ (5)
The motivation for the above metrics is that $\mathcal{Y}$ represents the ground truth, while $\hat{\mathcal{Y}}$ represents the prediction. However, in Figure 2, the two sets of emotions are considered equally important. Therefore, we use the Jaccard similarity coefficient to measure the similarity. This metric evaluates the similarity between two sets by comparing the size of their intersection to the size of their union:
$\mbox{Similarity}_{\mbox{s}}=\frac{|\mathcal{Y}\cap\hat{\mathcal{Y}}|}{|\mathcal{Y}\cup\hat{\mathcal{Y}}|}.$ (6)
Appendix I CLUE-Multi Analysis
In this section, we further analyze the reliability and comprehensiveness of CLUE-Multi from three aspects: discrete emotion recognition, dimensional emotion recognition, and visual clue statistics. Table 9 provides prompts and models for each part of the analysis.
Table 9: Prompts and corresponding models used in CLUE-Multi analysis.
Function (Model) Prompt
#1 Discrete Emotion Recognition (GPT-3.5) Please assume the role of an expert in the emotional domain. We provide clues that may be related to the emotions of the character. Based on the provided clues, identify the emotional states of the main characters. We provide a set of emotional candidates, please rank them in order of likelihood from high to low. The candidate set is {happy, angry, worried, sad, surprise, neutral}.
#2 Valence Estimation (GPT-3.5) As an expert in the emotional domain, we provide clues that may be related to the emotions of characters. Based on the provided clues, please identify the overall positive or negative emotional polarity of the main characters. The output should be a floating-point number ranging from -5 to +5. Here, -5 indicates extremely negative emotions, 0 indicates neutral emotions, and +5 indicates extremely positive emotions. Larger numbers indicate more positive emotions, while smaller numbers indicate more negative emotions. Please provide your judgment as a floating-point number with two decimal places, directly outputting the numerical result without including the analysis process.
#3 Visual Clue Analysis (GPT-3.5) Please assume the role of an expert in the field of emotions. We provide clues related to the emotions of the characters in the video. Please output the facial movements and body gestures involved in the description, separated by commas. The output format should be in list form.
Discrete Emotion Recognition.
Our dataset is based on MER2023, which provides relatively reliable one-hot labels. Therefore, we attempt to determine whether these one-hot labels can be identified from CLUE-Multi. This part of the analysis aims to verify whether CLUE-Multi can cover the traditional one-hot emotion recognition task. Experimental results indicate that the top-1 and top-2 scores can reach 93.48 and 96.89, respectively. Further analysis shows that the prediction errors are primarily due to the limitations of one-hot labels. For example, in Figure 1, the character shows a compound emotional state, including surprised, nervous, and unsatisfied. However, when we rank the candidate emotions, the output is: angry, surprised, worried, neutral, sad, happy. The top-1 label is angry, which differs from surprise in MER2023, leading to a prediction error. These results reveal the limitations of traditional one-hot labels in describing emotions.
Valence Estimation.
Besides discrete labels, MER2023 also provides relatively reliable valence scores. Therefore, we attempt to verify whether CLUE-Multi can used for valence estimation. Through experimental analysis, we observe that the PCC score between predictions and annotations can reach 0.88, indicating that CLUE-Multi also contains clues for dimensional emotion recognition.
Visual Clue Analysis.
Following that, we attempt to analyze the diversity of visual clues in CLUE-Multi. Through experimental analysis, we observe that each sample has an average of 4.95 visual clues. Therefore, we conclude that CLUE-Multi contains a wealth of clues that can help address discrete emotion recognition and valence estimation. Additionally, these results validate the completeness and reliability of CLUE-Multi.
Appendix J Visualization of CLUE-M/A/T/V
Figure 17 provides an example and visualizes CLUE-M/A/T/V.
Figure 17: Visualization of CLUE-M/A/T/V.
Appendix K Details of CLUE-MLLM
CLUE-MLLM directly utilizes the output from MLLM without any manual checking process. Table 10 provides model cards for different MLLMs. For each MLLM, we provide two types of prompts (see Table 11): one that ignores text and another that considers text. To ensure a fair comparison, we use similar prompts for audio, video, and audio-video LLMs.
Table 10: Model cards for MLLMs.
Model Link
SECap [49] https://github.com/thuhcsi/SECap
SALMONN [9] https://github.com/bytedance/SALMONN
Qwen-Audio [50] https://github.com/QwenLM/Qwen-Audio
Otter [51] https://github.com/Luodian/Otter
OneLLM [52] https://github.com/csuhan/OneLLM
PandaGPT [53] https://github.com/yxuansu/PandaGPT
VideoChat [54] https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat
VideoChat2 [55] https://github.com/OpenGVLab/Ask-Anything/tree/main/video_chat2
Video-LLaMA [56] https://github.com/DAMO-NLP-SG/Video-LLaMA
Video-LLaVA [57] https://github.com/PKU-YuanGroup/Video-LLaVA
Video-ChatGPT [58] https://github.com/mbzuai-oryx/Video-ChatGPT
LLaMA-VID [59] https://github.com/dvlab-research/LLaMA-VID
mPLUG-Owl [60] https://github.com/X-PLUG/mPLUG-Owl
Chat-UniVi [61] https://github.com/PKU-YuanGroup/Chat-UniVi
GPT-4V [62] https://openai.com/
Table 11: Prompts for extracting emotion-related descriptions using MLLMs.
Model Text Prompt
Audio LLM w/o As an expert in the field of emotions, please focus on the acoustic information in the audio to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual.
w/ Subtitle content of the audio: {subtitle}; As an expert in the field of emotions, please focus on the acoustic information and subtitle content in the audio to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the audio.
Video LLM w/o As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
w/ Subtitle content of the video: {subtitle}; As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual.
Audio-Video LLM w/o As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
w/ Subtitle content of the video: {subtitle}; As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
This paper tests different CLUE-MLLM generation strategies: S0, S1, and S2. Experimental results are shown in Table 12. We observe that S2 generally outperforms both S0 and S1. Therefore, we adopt S2 as the default strategy.
Table 12: Performance comparison of different strategies for generating CLUE-MLLM.
Model Strategy English Chinese
$\mbox{F}_{\mbox{s}}$ $\mbox{Precision}_{\mbox{s}}$ $\mbox{Recall}_{\mbox{s}}$ $\mbox{F}_{\mbox{s}}$ $\mbox{Precision}_{\mbox{s}}$ $\mbox{Recall}_{\mbox{s}}$
S0 34.75±0.02 40.41±0.03 30.48±0.01 31.08±0.10 35.71±0.15 27.51±0.07
Otter S1 22.54±0.05 26.05±0.08 19.86±0.04 25.06±0.04 29.14±0.03 21.99±0.05
S2 43.51±0.09 50.71±0.10 38.09±0.09 46.22±0.01 52.65±0.16 41.18±0.08
S0 26.99±0.01 29.18±0.08 25.10±0.04 28.70±0.01 30.95±0.00 26.76±0.03
PandaGPT S1 34.75±0.21 36.77±0.30 32.94±0.14 34.74±0.17 37.27±0.15 32.53±0.18
S2 45.89±0.20 50.03±0.01 42.38±0.33 47.33±0.04 53.01±0.08 42.75±0.11
S0 34.77±0.04 37.66±0.13 32.30±0.03 37.62±0.16 40.33±0.05 35.25±0.25
Video-ChatGPT S1 41.74±0.24 45.59±0.24 38.49±0.23 40.81±0.03 45.07±0.00 37.28±0.05
S2 50.52±0.06 54.03±0.04 47.44±0.07 54.73±0.00 61.15±0.10 49.52±0.06
S0 28.17±0.26 28.64±0.36 27.72±0.18 30.70±0.11 30.09±0.14 31.34±0.08
Video-LLaMA S1 34.43±0.16 35.82±0.20 33.15±0.11 34.01±0.25 35.16±0.22 32.94±0.26
S2 44.73±0.14 44.14±0.13 45.34±0.15 47.26±0.03 47.98±0.07 46.56±0.01
S0 31.95±0.02 31.73±0.13 32.17±0.10 34.53±0.02 33.53±0.01 35.60±0.05
VideoChat S1 45.10±0.07 46.24±0.05 44.01±0.10 44.25±0.09 44.76±0.02 43.75±0.16
S2 45.53±0.11 42.90±0.27 48.49±0.10 45.57±0.03 47.20±0.12 44.05±0.05
S0 35.70±0.06 43.08±0.00 30.47±0.09 35.27±0.01 41.16±0.00 30.86±0.01
VideoChat2 S1 37.56±0.07 44.62±0.00 32.43±0.10 38.71±0.10 45.14±0.13 33.88±0.08
S2 49.07±0.26 54.72±0.41 44.47±0.15 48.86±0.05 57.12±0.08 42.68±0.04
S0 39.21±0.14 40.56±0.15 37.94±0.12 40.53±0.33 40.44±0.24 40.62±0.43
mPLUG-Owl S1 45.80±0.06 47.49±0.04 44.22±0.07 47.97±0.04 49.33±0.03 46.69±0.05
S2 52.73±0.13 54.54±0.13 51.04±0.13 50.95±0.06 56.40±0.11 46.47±0.18
S0 40.71±0.10 41.38±0.25 40.07±0.04 43.45±0.23 43.24±0.30 43.66±0.16
SALMONN S1 39.79±0.03 39.54±0.01 40.05±0.06 41.43±0.13 41.11±0.03 41.76±0.22
S2 47.96±0.04 50.20±0.04 45.92±0.04 48.24±0.03 52.24±0.00 44.82±0.05
S0 30.64±0.06 41.92±0.00 24.14±0.08 30.50±0.05 40.84±0.13 24.33±0.03
Qwen-Audio S1 35.23±0.10 46.69±0.15 28.29±0.08 44.09±0.00 58.08±0.00 35.53±0.00
S2 38.13±0.05 49.42±0.18 31.04±0.00 41.14±0.07 53.71±0.00 33.34±0.09
S0 32.64±0.03 33.31±0.01 32.00±0.05 32.76±0.03 33.19±0.06 32.33±0.00
Video-LLaVA S1 30.19±0.02 34.10±0.03 27.08±0.05 31.93±0.11 33.40±0.19 30.58±0.04
S2 47.07±0.16 48.58±0.02 45.66±0.29 49.21±0.06 53.95±0.03 45.23±0.13
S0 35.14±0.14 36.71±0.15 33.69±0.14 33.30±0.04 33.12±0.06 33.48±0.03
LLaMA-VID S1 42.37±0.03 43.97±0.04 40.89±0.03 42.56±0.08 43.28±0.11 41.86±0.04
S2 51.25±0.09 52.71±0.18 49.87±0.00 52.01±0.02 57.30±0.00 47.61±0.03
S0 39.89±0.18 42.32±0.21 37.72±0.15 36.83±0.30 37.74±0.27 35.96±0.33
Chat-UniVi S1 47.94±0.19 50.96±0.20 45.26±0.18 47.02±0.00 48.07±0.00 46.01±0.00
S2 53.08±0.01 53.68±0.00 52.50±0.02 53.86±0.02 58.54±0.01 49.86±0.03
Appendix L GPT-based vs. Matching-based Metrics
Table 13 provides raw scores for GPT- and matching-based metrics. See Section 5 for more analysis.
Table 13: GPT-based vs. matching-based metrics. “$\mbox{P}_{\mbox{s}}$”, “$\mbox{R}_{\mbox{s}}$”, “B1”, “B4”, “’M’, and “Rl” are abbreviations for $\mbox{Precision}_{\mbox{s}}$, $\mbox{Recall}_{\mbox{s}}$, BLEU1, BLEU4, METEOR, and ROUGEl, respectively.
MLLM L V A English Chinese
GPT-based Matching-based GPT-based Matching-based
$\mbox{F}_{\mbox{s}}$ $\mbox{P}_{\mbox{s}}$ $\mbox{R}_{\mbox{s}}$ B1 B4 M Rl $\mbox{F}_{\mbox{s}}$ $\mbox{P}_{\mbox{s}}$ $\mbox{R}_{\mbox{s}}$ B1 B4 M Rl
Qwen-Audio $\surd$ $\times$ $\surd$ 38.13 49.42 31.04 21.87 06.55 21.65 20.81 41.14 53.71 33.34 27.64 12.07 26.09 25.24
OneLLM $\surd$ $\times$ $\surd$ 42.84 45.92 40.15 33.81 08.54 28.00 22.46 46.17 52.07 41.47 42.75 16.60 34.42 26.81
Otter $\surd$ $\surd$ $\times$ 43.51 50.71 38.09 27.26 07.55 23.42 21.05 46.22 52.65 41.18 35.35 14.41 29.34 25.91
Video-LLaMA $\surd$ $\surd$ $\times$ 44.73 44.14 45.34 28.76 06.41 31.22 20.41 47.26 47.98 46.56 34.88 12.13 37.61 24.25
VideoChat $\surd$ $\surd$ $\times$ 45.53 42.90 48.49 26.44 05.41 30.58 19.11 45.57 47.20 44.05 31.36 10.86 37.48 22.57
PandaGPT $\surd$ $\surd$ $\surd$ 45.89 50.03 42.38 33.69 07.64 30.29 22.07 47.33 53.01 42.75 43.02 15.83 37.94 26.87
Video-LLaVA $\surd$ $\surd$ $\times$ 47.07 48.58 45.66 33.48 08.25 29.68 22.34 49.21 53.95 45.23 42.72 15.97 36.87 26.90
SALMONN $\surd$ $\times$ $\surd$ 47.96 50.20 45.92 31.89 07.19 28.42 20.99 48.24 52.24 44.82 39.00 14.00 35.12 25.35
VideoChat2 $\surd$ $\surd$ $\times$ 49.07 54.72 44.47 31.60 08.10 26.61 21.65 48.86 57.12 42.68 41.18 16.15 33.54 26.80
Video-ChatGPT $\surd$ $\surd$ $\times$ 50.52 54.03 47.44 32.64 07.65 30.25 22.01 54.73 61.15 49.52 41.96 15.50 38.18 26.35
OneLLM $\surd$ $\surd$ $\times$ 50.52 55.93 46.06 32.19 08.10 28.44 22.25 51.44 56.43 47.26 41.31 15.15 35.15 25.98
LLaMA-VID $\surd$ $\surd$ $\times$ 51.25 52.71 49.87 33.81 08.26 30.31 22.36 52.01 57.30 47.61 43.01 16.23 37.92 27.20
mPLUG-Owl $\surd$ $\surd$ $\times$ 52.73 54.54 51.04 33.04 07.75 30.24 21.75 50.95 56.40 46.47 41.69 15.16 37.81 26.39
Chat-UniVi $\surd$ $\surd$ $\times$ 53.08 53.68 52.50 32.80 07.83 31.12 22.15 53.86 58.54 49.86 40.76 15.05 38.75 26.43
GPT-4V $\surd$ $\surd$ $\times$ 55.51 48.52 64.86 39.40 18.41 43.67 32.60 57.21 54.61 60.07 45.45 29.08 53.76 40.37
Appendix M Emotion Wheel
[…]
(a) W1
(b) W2
(c) W3
(d) W4
(e) W5
Figure 18: Emotion wheels. This paper selects five representative emotion wheels (please zoom in to clearly view the emotional hierarchy): (a) W1 (b) W2 (c) W3 (d) W4 (e) W5